OpenStudy (anonymous):

Which statement is true for the expression a + (3 + 6y)?

OpenStudy (anonymous):

A) It can be written as (a + 3) + 6y by using associative property of addition. B) It can be written as a + 3 + 6y + (3 + 6y) by using distributive property of addition. C) It can be written as a + 3 + 6y + (3 + 6y) by using associative property of addition. D) It can be written as (a + 3) + 6y by using distributive property of addition.

OpenStudy (ghazi):

associative law says (a+b)+c= a+(b+c)

OpenStudy (ghazi):

distributive says a(b+c)= ab+bc

OpenStudy (anonymous):

.. I still don't really understand. Is the answer A?

OpenStudy (anonymous):

I Believe the answer is A. Im Not So Sure But You Should Google It. c:

OpenStudy (anonymous):

Ohkiee c: Thanks!

OpenStudy (ghazi):

yes answer is A